OpenStudy (anonymous):

7) Find S25 for -7+ -2 + 3 + 8 +…

OpenStudy (irishboy123):

hint: s1 = -7 s2 = -7 + (1)5 s3 = -7 + 2(5) ........... you can keep going to 25(!) or find the formula and use it

OpenStudy (anonymous):

123?

OpenStudy (anonymous):

@IrishBoy123